OpenStudy (anonymous):

Solve: 125x^3 + 343=0 Find all Complex Roots.

OpenStudy (anonymous):

try http://www.freemathhelp.com/ it help with almost everything

OpenStudy (campbell_st):

this is the sum of 2 cubics (5x)^3 and (7)^3 then it can be factorised to (5x + 7)(25x^2 - 35x + 49) the quadratic needs to be solved by general quadratic formula \[x = (35 \pm \sqrt{245})/10\] then the solutions are \[x = (35 \pm 7\sqrt{5})/10\] and x = - 7/5

OpenStudy (anonymous):

so could that be -7/5, and 35 plus/minus 35i sqrt 3 over 50

OpenStudy (campbell_st):

yep... there are 3 roots